OpenStudy (anonymous):

A car brakes and decelerates uniformly. After t seconds, it has travelled a distance of d metres where d is related to t by the formula, d= 40t-2t^2, t is greater or equal to 10. How far does the car travel in the next second?

OpenStudy (anonymous):

Please break it down for me thanks.

OpenStudy (anonymous):

i got d=36t+38-2t^2 but the answer is (38-4t) . Please explain why is this so.

OpenStudy (jack1):

ok this equation comes from a physics formula you'll come across: s = ut + 1/2 a t^2 s = distance u = initial speed a = acceleration / deceleration t = time so from this we know: car's initial speed = 40 meters/sec cars deceleration = -4 meters/sec^2 d= 40t - 2t^2 \[d= 40t-2t^2\] so if t = 10 d = 400 - 200 = 200 in the next second (t = 11) distance = d= 40t-2t^2 d = 440 - 242 d = 198 Col 1 = Total distance travelled col 2 = time (seconds) col 3 = distance travelled since the last second 38 1 38 72 2 34 102 3 30 128 4 26 150 5 22 168 6 18 182 7 14 192 8 10 198 9 6 200 10 2 198 11 -2 192 12 -6 182 13 -10 168 14 -14 150 15 -18 128 16 -22 102 17 -26 72 18 -30 38 19 -34 0 20 -38

OpenStudy (jack1):

so from time = 10 seconds and onwards, it seems like the car is going backwards

OpenStudy (jack1):

as between t = 10 and t = 11, the car travels - 2 m compared to the last second